The sum of two numbers is 35. The smaller number is 5 less than the larger number. What are the numbers?
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

The sum of two numbers is 35. The smaller number is 5 less than the larger number. What are the numbers?
============
x + y = 35
x - y = 5

Answer by ikleyn(52810)

.
Hey,
you can solve it mentally, very easy.

Diminish the larger number by 5 units (mentally).

Then you will have two equal numbers whose sum is 30.

Hence . . . what?

The smaller number is 30/2 = 15. The larger number is 15+5 = 20.

That's all.


There is no need to use any equations.